# Multimodal-enhanced Federated Recommendation: A Group-wise Fusion Approach

### v1

@@ 5.6. Privacy Protection-Enhanced GFMFR
To further enhance the privacy protection of the proposed model, we integrate local differential privacy (LDP) by injecting Laplacian noise into client-side model parameters. Specifically, we experiment with five levels of noise intensity, setting the noise scale to $\delta=\{1,2,3,4,5\}$, approximating LDP settings under unit sensitivity, with injected noise levels corresponding to decreasing the privacy budget values in the range $[1,0.2]$. This setup allows us to analyze the trade-off between privacy and performance under different LDP configurations. As shown in Table 2, increasing the noise intensity improves privacy protection but leads to a gradual decline in model accuracy. In practical deployments, a moderate noise level can serve as a balanced trade-off, providing sufficient privacy without significantly sacrificing recommendation quality.
Table 2. Performance of integrating LDP into our proposed GFMFR under different noise scales. “H@50” and “N@50” are abbreviations for HR@50 and NDCG@50.
Dataset $\delta=0$ $\delta=1$ $\delta=2$ $\delta=3$ $\delta=4$ $\delta=5$
H@50 N@50 H@50 N@50 H@50 N@50 H@50 N@50 H@50 N@50 H@50 N@50
Beauty 0.2024 0.0730 0.1981 0.0633 0.1798 0.0591 0.1726 0.0587 0.1631 0.0496 0.1566 0.0443
Tools_and_Home 0.2269 0.0612 0.2025 0.0538 0.1708 0.0442 0.1660 0.0394 0.1550 0.0371 0.1529 0.0334
Toys_and_Games 0.2979 0.2613 0.2897 0.2471 0.2861 0.2402 0.2841 0.2393 0.2814 0.2385 0.2775 0.2365
Digital_Music 0.2943 0.2820 0.2786 0.2594 0.2668 0.2489 0.2650 0.2431 0.2603 0.2394 0.2551 0.2263
Office_Products 0.0614 0.0198 0.0119 0.0031 0.0109 0.0027 0.0105 0.0026 0.0100 0.0025 0.0095 0.0023
@@ 6. Conclusion
To address the limitations of existing federated recommender systems in handling rich multimodal content, we propose GFMFR, a flexible and efficient framework that systematically integrates multimodal information into the FR pipeline. By offloading the heavy multimodal representation learning to the server side and equipping clients with compact, distilled representations, GFMFR effectively alleviates client-side computation and communication burdens. The introduction of group-aware multimodal aggregation allows for adaptive knowledge sharing among users with similar preferences, while the preference-guided distillation ensures that multimodal signals are aligned with personalized recommendation objectives. Extensive empirical studies demonstrate the superiority of GFMFR in both performance and system efficiency across multiple datasets.

### v2

@@ 5.6. Privacy Protection-Enhanced GFMFR
To further enhance the privacy protection of the proposed model, we integrate local differential privacy (LDP) by perturbing all shared parameters with Laplacian noise prior to transmission to the server. Specifically, we experiment with five levels of noise intensity by setting the noise scale $\delta\in\{1,2,3,4,5\}$ under unit sensitivity, which corresponds to privacy budgets $\epsilon\in\{1.0,0.5,0.33,0.25,0.2\}$. As shown in Table 2, increasing the noise intensity improves privacy protection but leads to a gradual decline in model accuracy. In practical deployments, a moderate noise level can serve as a balanced trade-off, providing sufficient privacy without significantly sacrificing recommendation quality.
Table 2. Performance of integrating LDP into our proposed GFMFR under different noise scales $\delta$.
Dataset Noise scale $\bm{\delta}$=0 $\bm{\delta}$=1 $\bm{\delta}$=2 $\bm{\delta}$=3 $\bm{\delta}$=4 $\bm{\delta}$=5
Beauty HR@50 $0.2024$ $0.1981$ $0.1798$ $0.1726$ $0.1631$ $0.1566$
NDCG@50 $0.0730$ $0.0633$ $0.0591$ $0.0587$ $0.0496$ $0.0443$
TH HR@50 $0.2269$ $0.2025$ $0.1708$ $0.1660$ $0.1550$ $0.1529$
NDCG@50 $0.0612$ $0.0538$ $0.0442$ $0.0394$ $0.0371$ $0.0334$
TG HR@50 $0.2979$ $0.2897$ $0.2861$ $0.2841$ $0.2814$ $0.2775$
NDCG@50 $0.2613$ $0.2471$ $0.2402$ $0.2393$ $0.2385$ $0.2365$
DM HR@50 $0.2943$ $0.2786$ $0.2668$ $0.2650$ $0.2603$ $0.2551$
NDCG@50 $0.2820$ $0.2594$ $0.2489$ $0.2431$ $0.2394$ $0.2263$
OP HR@50 $0.0614$ $0.0119$ $0.0109$ $0.0105$ $0.0100$ $0.0095$
NDCG@50 $0.0198$ $0.0031$ $0.0027$ $0.0026$ $0.0025$ $0.0023$
@@ 6. Conclusion
To address the limitations of existing federated recommender systems in handling rich multimodal content, we propose GFMFR, a flexible and efficient framework that systematically integrates multimodal information into the FR pipeline. By offloading the heavy multimodal representation learning to the server side and equipping clients with compact, distilled representations, GFMFR effectively alleviates client-side computation and communication burdens. The introduction of group-aware multimodal aggregation allows for adaptive knowledge sharing among users with similar preferences, while the preference-guided distillation ensures that multimodal signals are aligned with personalized recommendation objectives. Extensive empirical studies demonstrate the consistent superiority of GFMFR in both recommendation performance and system efficiency across multiple datasets.
Acknowledgements.
Chunxu Zhang, Weipeng Zhang, Zhiheng Xue and Bo Yang are supported by the National Natural Science Foundation of China under Grant Nos. U22A2098, 62172185, 62206105 and 62202200; the Major Science and Technology Development Plan of Jilin Province under Grant No.20240212003GX, the Major Science and Technology Development Plan of Changchun under Grant No.2024WX05. Riting Xia is supported by the National Natural Science Foundation of China under Grant Nos. 62506177; the Inner Mongolia Autonomous Region Natural Science Foundation under Grant No.2025QN06010.
Appendix A Experimental Setup
A.1. Dataset Summary
These datasets, spanning diverse domains, sparsity levels, and scales, provide users’ interaction records as well as item metadata (e.g., titles and images) that constitute the raw multimodal signals for our study. Following the data preprocessing strategy in (Zhang et al., 2023a), we exclude users and items with fewer than 10 interactions. For the dataset split, we adopt the prevalent leave-one-out evaluation setting (He et al., 2017; Zhang et al., 2023a; Li et al., 2024b). The statistics are summarized in Table 4.
A.2. Baseline Details
We first introduce the three representative ID-based FR backbone models. Specifically, FedNCF is one of the most representative FR models, it deploys a neural collaborative filtering recommendation model (He et al., 2017) on each client and learns the shared item embedding and prediction function for all clients. PFedRec(Zhang et al., 2023a) and FedRAP(Li et al., 2024b) are two state-of-the-art FR models that advocate learning personalized item embeddings for each client to better capture individual user preferences. Then, we give a brief summary of three existing multimodal-based FR baselines. FedMMR (Li et al., 2024a) introduces multimodal information into FR by incorporating a server-side step that aligns various modality representations with global item embeddings. AMMFRS (Feng et al., 2024) and FedMR (Li et al., 2024c) enhance local item representation by integrating multimodal features, using self-attention and mixture-of-experts mechanisms, respectively.
A.3. Implementation details
Specifically, we predict over the full item set and calculate the metrics based on the rank of the test item within the top-50 recommendation list. We report the average results over five independent runs. All experiments are implemented in PyTorch and conducted on a machine with three NVIDIA GeForce RTX 3090 GPUs. For multimodal representation learning, we employ CLIP based on the Vision Transformer Base (ViT-B/32) architecture (Radford et al., 2021; Dosovitskiy et al., 2020). Missing visual features are imputed with the average values from available items, while missing titles are replaced with a placeholder string ”There is no title.” To ensure a fair comparison, we set the latent dimension to 512, the batch size to 256, and the number of communication rounds to 100, which we empirically find sufficient for convergence. We additionally conduct an exhaustive hyperparameter search for all baseline methods, grid-searching the learning rate over {0.1, 0.01, 0.001, 0.0001} and the weight decay over {1e-3, 1e-4, 1e-5}. For other method-specific hyperparameters, we either retain the values from their original publications or optimize them on our validation set until performance plateaued. Consistent with prior work (He et al., 2017; Zhang et al., 2023a), we adopt a negative sampling strategy with four negative instances per positive instance during training.
Table 4. Dataset statistics.
Dataset #Users #Items #Interactions Sparsity
Beauty 253 356 2,535 97.19%
Tools_and_Home 242 398 3,606 96.26%
Toys_and_Games 1,146 1,403 23,820 98.52%
Digital_Music 1,517 2,306 35,395 98.99%
Office_Products 6,338 4,461 97,227 99.65%
Appendix B Additional Experiments
B.1. Efficiency Comparison Results
In real-world deployments of FR systems, efficiency is as critical as performance, particularly given the limited storage and computation resources on client devices and the frequent communication required between clients and the server during federated optimization. The efficiency comparison results are summarized in Table 3. The Client_Storage_Overhead column includes two metrics: (1) the size of multimodal representations stored on the client, and (2) the parameter count of the local recommendation model. The Communication_Overhead column reports: (1) the initial distribution cost of the multimodal representations, and (2) the per-round transmission cost between clients and the server. The Client_Epoch_Time column denotes the per-epoch training time of the client-side recommendation model.
For baselines AMMFRS, FedMR, and FedMLP, storing item multimodal features locally and designing complex fusion modules increase client storage burden; FedMMR’s storage overhead stems from its architecture integrating an elaborate contrastive loss. The complexity of these local models lengthens client training and enlarges per-round transmission cost between clients and server. Additionly, AMMFRS, FedMR, and FedMLP need to transmit full initial multimodal representations to all clients before training, incurring substantial communication overhead. In contrast, our method keeps item multimodal content on the server and employs a high-capacity encoder to generate representations, alleviating client computational and storage burden; only low-dimensional group-preference vectors (e.g., 2 vs. 512) are exchanged, introducing minimal communication overhead. Consequently, our method demonstrates superior efficiency, making it a more practical solution for real-world deployment under resource constraints.
Figure 5. Hyper-parameter analysis results on dataset Beauty.
Figure 6. Hyper-parameter analysis results on dataset Tools_and_Home
Figure 7. Hyper-parameter analysis results on dataset Toys_and_Game.
Figure 8. Hyper-parameter analysis results on dataset Digital_Music
Figure 9. Hyper-parameter analysis results on dataset Office_Products
B.2. Further Hyper-Parameter Analysis Results
Number of user groups $\bm{g}$.
Based on the user scale of each dataset, we vary the number of user groups to investigate its impact on recommendation performance. Specifically, we test group numbers of $\{5,10,20,50,100\}$ for Beauty and Tools_and_Home, $\{5,10,20,100,500\}$ for Toys_and_Games and Digital_Music, and $\{5,50,100,500,1000\}$ for Office_Products. As shown in Figure 5, 6, 7, 8, 9 (a), experimental results across five datasets demonstrate that the number of user groups significantly influences performance, with the optimal value strongly depending on dataset-specific characteristics. This finding suggests that our method effectively adapts to the unique properties of each dataset by partitioning users into groups, thereby capturing diverse item preferences across these groups.
Number of local training epochs $\bm{H}$.
We evaluate the impact of the number of local training epochs from the set $\{1,3,5,10,20\}$ across all datasets. As shown in Figure 5, 6, 7, 8, 9 (b), experimental results indicate that setting the number of local epochs to 5 yields optimal or near-optimal performance in all cases. Too few local epochs result in underfitted and noisy updates, whereas an excessive number induces overfitting to local data and escalates computational costs, ultimately degrading model performance.
Client sampling ratio $\bm{\beta}$.
We examine the impact of the client sampling ratio across all datasets by testing values in $\{0.2,0.4,0.6,0.8,1.0\}$. As shown in Figure 5, 6, 7, 8, 9(c), different sampling ratios generally yield comparable final performance. While a lower sampling ratio involves fewer clients per communication round, this can be offset by increasing the number of rounds to achieve convergence. In contrast, higher sampling ratios accelerate convergence and improve computational efficiency.